OpenStudy (anonymous):

Solve: −2(x − 5) = −4 7 −7 negative 1 over 2 −3

OpenStudy (anonymous):

-2x+10=-4 -2x=-14 x=14/2=7

OpenStudy (anonymous):

saifoo - you forgot to multiply the -5 by -2

OpenStudy (saifoo.khan):

Dang! i better head to bed! :/

OpenStudy (anonymous):

Haha :D

OpenStudy (saifoo.khan):

It's 3:35am here! O_O